# Localised Natural Causal Learning Algorithms for Weak Consistency Conditions

### v1

@@ Abstract
By relaxing conditions for “natural” structure learning algorithms—a family of constraint-based algorithms containing all exact structure learning algorithms under the faithfulness assumption, we define localised natural structure learning algorithms (LoNS). We also provide a set of necessary and sufficient assumptions for consistency of LoNS, which can be thought of as a strict relaxation of the restricted faithfulness assumption. We provide a practical LoNS algorithm that runs in exponential time, which is then compared with related existing structure learning algorithms, namely PC/SGS and the relatively recent Sparsest Permutation algorithm. Simulation studies are also provided.
@@ 1 Introduction
Inferring causal relationships have always been of great interest in different fields, with some frameworks like potential outcomes and graphs gaining prominence amongst the causality community. A main goal of graph-based causal inference is causal discovery: given data, we would like to uncover the underlying causal structure in the form of a true causal graph, on which conventional graph-based causal inference techniques hinges. We will mostly be concerned with the setting of observational data only, which is a reasonable assumption when interventional data in the form of randomised control trials are unavailable, to which the true causal graph is recoverable up to its graphical separations. Current causal discovery approaches can generally be categorised into: score-based approaches [Chickering, 2002] and constraint-based approaches [Spirtes et al., 2000]. Here, we will mostly be concerned with the latter.
Assumptions are needed for constraint-based approaches, otherwise many causal structures representing the same data may be obtained, and results in making vacuous causal statements. Amongst these, the most common and widely known is faithfulness assumption, where every conditional independence in the data generating distribution is exactly represented by the true causal graph [Zhang and Spirtes, 2008]; most constraint-based learning approaches such as PC and SGS provably return the true causal graph up to its graphical separations. However in practice and theory, the condition is shown be too strong at times [Uhler et al., 2013].
Efforts to relax the faithfulness assumption include the Sparsest Permutation (SP) algorithm by Raskutti and Uhler [2018], which provably returns the graphical separations of the true causal graph under strictly weaker assumption than faithfulness, at the expense of factorial run time by permuting the causal variables. Greedy approaches to speed up the SP algorithm [Solus et al., 2021, Lam et al., 2022] have been proposed, however these algorithms only return the true causal graph under strictly stronger conditions than SP.
Addressing this, Sadeghi and Soo [2022] proposed the class of “natural” structure learning algorithms, which under the faithfulness assumption, encompasses constraint-based approaches such as SGS/PC algorithms. In addition, natural structure learning algorithms are also proven to return the true causal graph up to graphical separations under well defined assumptions that are shown to be strictly weaker than faithfulness. Thus, the objective of this paper is as follows: 1.) To further weaken the consistency conditions by defining a localised version of natural structure learning algorithms. 2.) To provide a practical algorithm of this type that works under these conditions.
@@ 2.1 Graphs
We first introduce the relevant concepts in graphical models, as well as some existing results in literature. In this work, unless noted otherwise, graphs will be implicitly assumed to be a directed acyclic graph (DAG) that is a graph over the set of nodes $V=\{1,...,n\}$, with directed edges such that there does not exist a sequence of directed edges from a node to itself. We denote $an_{G}(C)$ as the set of nodes $i\not\in C$ such that there exists a sequence of directed edges from $i$ to some $j\in C$ in $G$.
We denote $A\perp_{G}B\,|\,C$ as graphical separation (in the case of DAGs, this can be understood as d-separation) between $A,B$ given $C$, where $A,B,C\subseteq V$ is disjoint, in graph $G$. A set of random variables $X=(X_{1},...,X_{n})$ with joint distribution $P$ is associated to the set of nodes $V$. We denote $A\mbox{$\>\perp\perp$ }B\,|\,C$ as conditional independence of $X_{i},i\in A$ and $X_{j},j\in B$ given $X_{k},k\in C$. We relate the two notions together using Markov property:
@@
Distribution $P$ is Markovian to $G$ if we have: $A\perp_{G}B\,|\,C\Rightarrow A\mbox{$\>\perp\perp$ }B\,|\,C$ for all disjoint $A,B,C\subseteq V$.
@@
$P$ is faithful to $G$ if we have: $A\perp_{G}B\,|\,C\iff A\mbox{$\>\perp\perp$ }B\,|\,C$ for all disjoint $A,B,C\subseteq V$.
Sadeghi [2017] has shown that $P$ being faithful to DAG $G$ implies that $P$ satisfies ordered upward stability and ordered downward stability wrt $G$, defined in the case of DAGs, as follows:
@@
    (Ordered upward stability (OUS))
    $P$ satisfies ordered upward stability wrt $G$ if: for all $i,j,k$ and $C\subseteq V\backslash\{i,j,k\}$, such that $k\in an_{G}(i,j)$, the following holds: $i\mbox{$\>\perp\perp$ }j\,|\,C\Rightarrow i\mbox{$\>\perp\perp$ }j\,|\,C\cup k$.
@@
    (Ordered downward stability (ODS))
    $P$ satisfies ordered downward stability wrt $G$ if: for all $i,j,k$ and $C\subseteq V\backslash\{i,j,k\}$, such that $k\not\in an_{G}(i,j,C)$, the following holds: $i\mbox{$\>\perp\perp$ }j\,|\,C\cup k\Rightarrow i\mbox{$\>\perp\perp$ }j\,|\,C$.
@@
Denote $sk(G)$ to be the skeleton of graph $G$, formed by removing all arrowheads from edges in $G$. A v-configuration is a set of nodes $i,k,j$ such that $i$ and $j$ are connected to $k$, but $i$ and $j$ are not connected, and will be represented as $i\sim k\sim j$. A v-configuration oriented as $i\xrightarrow{}k\xleftarrow{}j$ is a collider, otherwise the v-configuration is a non-collider.
@@
To relate $sk(G)$ with distribution $P$, we define $sk(P)$ as:
Definition 4 ($sk(P)$).
Given a distribution $P$, $sk(P)$ is the undirected graph with node set $V$, such that for all $i,j\in V$, $i$ is adjacent to $j$ if and only if there does not exist any $C\subset V\backslash\{i,j\}$ such that $i\mbox{$\>\perp\perp$ }j\,|\,C$.
Note that $sk(P)$ defined is the output of the skeleton building step of SGS/PC algorithm under faithfulness. Then if $P$ is adjacency faithful wrt $G$, $sk(P)=sk(G)$.
@@ 2.2 Natural Structure Learning Algorithms
Let $P$ be Markovian to the true causal graph $G_{0}$, then a causal learning algorithm aims to recover the graph $G_{0}$, up to the MEC, in which case we say that the algorithm is consistent. To relax the faithfulness assumption, Sadeghi and Soo [2022] have introduced natural structure learning algorithms, defined as:
@@
$P$ is V-stable if:
for all v-configurations $i\sim k\sim j$ in $sk(P)$, and $C\subseteq V\backslash\{i,j,k\}$, $i\mbox{$\>\perp\perp$ }j\,|\,C$ and $i\mbox{$\>\perp\perp$ }j\,|\,C\cup k$ cannot both hold.
@@
This is a definition on $P$ itself, and is an implication of the well known singleton transitive axiom, under adjacency faithfulness.
Proposition 1 (Sadeghi and Soo [2022]).
$G(P)$ and $G_{0}$ are Markov equivalent if the following holds:
@@
In Sadeghi and Soo [2022], Condition 1 above is given in terms of converse pairwise Markovian instead of adjacency faithfulness, this is due to attempts in characterising the consistency conditions in terms of Structural Causal Models (SCM). However, only the weaker adjacency faithfulness is needed and here we are focused on relaxing conditions.
By Example 21 in Sadeghi and Soo [2022], it can be seen that combined, these conditions are strictly weaker than restricted faithfulness, which is the weakest known consistency condition for SGS/PC [Raskutti and Uhler, 2018].
Under the faithfulness assumption, it is shown that constraint-based structure learning algorithms are natural. However it is unclear whether these algorithms are still natural structure learning algorithms once the faithfulness assumption is relaxed. Thus, without assuming faithfulness, we aim to provide a general natural structure learning algorithm that relaxes the consistency conditions in Proposition 1.
As per usual in constraint-based causal learning, we assume the availability of a conditional independence oracle: given a probability distribution $P$, we can determine with certainty whether conditional independence statements are true. In practice, conditional independence statements need to be estimated from the data using methods such as HSIC testing [Gretton et al., 2007], and is shown to be in general, a hard problem [Shah and Peters, 2020].
@@ 3.1 Theory
Definition 8 (V-OUS and Collider-Stable).
Distribution $P$ is V-OUS and collider-stable wrt DAG $G$ if for all v-configuration $i\sim k\sim j$ in $G$:
@@
    (V-Ordered upward stability (V-OUS))
    If $i\sim k\sim j$ is a non-collider; for all $C\subseteq V\backslash\{i,j,k\}$, $i\mbox{$\>\perp\perp$ }j\,|\,C\Rightarrow i\mbox{$\>\perp\perp$ }j\,|\,C\cup k$.
@@
    (Collider-stability)
    If $i\xrightarrow[]{}k\xleftarrow{}j$; there exists $C^{\prime}\subseteq V\backslash\{i,j,k\}$, such that $i\mbox{$\>\perp\perp$ }j\,|\,C^{\prime}$.
Collider-stable has related notions such as orientation faithfulness, which states that the graph is faithful up to v-configurations in the graph [Raskutti and Uhler, 2018]. However collider-stable is much weaker, even than the Markovian assumption.
@@
V-OUS and collider-stable can be seen as local versions of ordered stabilities for the purposes of learning DAGs. In the case of DAGs, V-OUS can be seen as a relaxation of ous since the if-then statement only has to hold for $i,j,k$ that are non-colliders in $G$. Likewise, collider-stable is implied by ods and can be seen as a relaxation.
@@
Note that the above is the same with natural structure learning algorithms, just that one of the requirements is relaxed, namely ordered stabilities to V-OUS and collider-stable, thus just like natural structure learning algorithms, all constraint-based algorithms that work under faithfulness are localised natural.
@@
A V-OUS and Collider-Stable orientation rule wrt $P$ is defined as an assignment of v-configurations $i\sim k\sim j$ in $sk(P)$ into colliders and non-colliders as follows:
@@
    If for all $C$ such that $i\mbox{$\>\perp\perp$ }j\,|\,C$, $k\in C$, then assign $i\sim k\sim j$ to be a non-collider.
@@ A DAG $G$ is said to satisfy the V-OUS and collider-stable orientation rule wrt $P$, if $G$ satisfies:
    For all v-configurations $i\sim k\sim j$ in $G$, the following holds: via the orientation rule in Definition 10:
@@
We then have the following characterisation:
@@
We can now apply the V-OUS and collider-stable orientation rule wrt $P$ to assign the v-configurations in $sk(P)$. Note that the assignment may be incomplete, in the sense that some v-configurations may not satisfy either of the conditions in Definition 10 and are therefore unassigned. Modified V-stability is then defined as when this ambiguity can be resolved using the constraint that the graph is a DAG, as illustrated in Figure 1.
@@
Distribution $P$ is modified V-stable, if the v-configurations of DAGs that satisfy the V-OUS and collider-stable orientation rule wrt $P$ is unique.
Figure 1: Different assigned $sk(P)$. Label $\mathcal{N,C,U}$ denotes the v-configuration that is assigned to be a non-collider, collider and unassigned, respectively. In a), the unassigned v-configuration $4\sim 3\sim 2$ is constrained to be a collider due to acylicity of a DAG. In b), from lack of bidirectedness in a DAG, unassigned v-configuration $1\sim 2\sim 3$ is constrained to be a non-collider. In c), since there exists DAGs such that v-configuration $2\sim 3\sim 4$ can be either a collider or a non-collider, this ambiguity cannot be resolved. We say that $P$ that produces assigned $sk(P)$ whose unassigned v-configurations can be resolved using DAG constraints (for example a,b) is modified V-stable.
Under V-stability, all v-configurations in $sk(P)$ must satisfy either of the conditions in Definition 10, leaving no v-configurations in $sk(P)$ unassigned, thus V-stability implies modified V-stability.
@@
We then have the following, for the true causal graph $G_{0}$:

### v2

@@ Abstract
By relaxing conditions for “natural” structure learning algorithms, a family of constraint-based algorithms containing all exact structure learning algorithms under the faithfulness assumption, we define localised natural structure learning algorithms (LoNS). We also provide a set of necessary and sufficient assumptions for consistency of LoNS, which can be thought of as a strict relaxation of the restricted faithfulness assumption. We provide a practical LoNS algorithm that runs in exponential time, which is then compared with related existing structure learning algorithms, namely PC/SGS and the relatively recent Sparsest Permutation algorithm. Simulation studies are also provided.
@@ 1 Introduction
Inferring causal relationships has always been of great interest in different fields, with some frameworks like potential outcomes and graphs gaining prominence amongst the causality community. A main goal of graph-based causal inference is causal discovery; given data, we would like to uncover the underlying causal structure in the form of a true causal graph, on which conventional graph-based causal inference techniques hinge. We will mostly be concerned with the setting of observational data only, such as when interventional data in the form of randomised control trials are unavailable, from which the true causal graph is recoverable up to its graphical separations. Current causal discovery approaches can generally be categorised into score-based approaches [score] and constraint-based approaches [const]. Here, we will mostly be concerned with the latter.
Assumptions are needed for constraint-based approaches, otherwise many causal structures representing the same data may be obtained, resulting in vacuous causal statements. Amongst these, the most common and widely known is the faithfulness assumption, where every conditional independence in the data generating distribution is exactly represented by the true causal graph [Faith]; most constraint-based learning approaches such as PC and SGS provably return the true causal graph up to its graphical separations. However in practice and theory, the condition can be too strong at times [faithbad].
Efforts to relax the faithfulness assumption include the Sparsest Permutation (SP) algorithm by UhlSP, which provably returns the graphical separations of the true causal graph under strictly weaker assumption than faithfulness, at the expense of factorial run time by permuting the causal variables. Greedy approaches to speed up the SP algorithm [UhlGSP, lam] have been proposed, however these algorithms only return the true causal graph under strictly stronger conditions than SP.
Addressing this, [Sad] proposed the class of “natural” structure learning algorithms, which under the faithfulness assumption, encompasses constraint-based approaches such as SGS/PC algorithms. In addition, natural structure learning algorithms are also proven to return the true causal graph up to graphical separations under well defined assumptions that are shown to be strictly weaker than faithfulness. Thus, the objective of this paper is as follows: 1.) To further weaken the consistency conditions by defining a localised version of natural structure learning algorithms. 2.) To provide a practical algorithm of this type that works under these conditions.
@@ 2.1 Graphs
[…]
We denote $A\perp_{G}B\,|\,C$ as graphical separation in graph $G$ between $A,B$ given $C$, where $A,B,C\subseteq V$ are disjoint; in the case of DAGs, this can be understood as d-separation. A set of random variables $X=(X_{1},...,X_{n})$ with joint distribution $P$ is associated to the set of nodes $V$. We denote $A\mbox{$\>\perp\perp$ }B\,|\,C$ as conditional independence of $(X_{i})_{i\in A}$ and $(X_{j})_{j\in B}$ given $(X_{k})_{k\in C}$. We relate the two notions together using Markov property:
@@
A distribution $P$ is Markovian to $G$ if $A\perp_{G}B\,|\,C\Rightarrow A\mbox{$\>\perp\perp$ }B\,|\,C$ for all disjoint $A,B,C\subseteq V$.
@@
A distribution $P$ is faithful to $G$ if $A\perp_{G}B\,|\,C\iff A\mbox{$\>\perp\perp$ }B\,|\,C$ for all disjoint $A,B,C\subseteq V$.
orderfaith has shown that $P$ being faithful to DAG $G$ implies that $P$ satisfies ordered upward stability and ordered downward stability wrt $G$, defined in the case of DAGs, as follows:
@@
    (Ordered upward stability (OUS)). $P$ satisfies ordered upward stability wrt $G$ if for all $i,j,k$ and $C\subseteq V\backslash\{i,j,k\}$, such that $k\in\text{\emph{an}}_{G}(i,j)$, we have $i\mbox{$\>\perp\perp$ }j\,|\,C\Rightarrow i\mbox{$\>\perp\perp$ }j\,|\,C\cup\{k\}$.
@@
    (Ordered downward stability (ODS)). $P$ satisfies ordered downward stability wrt $G$ if for all $i,j,k$ and $C\subseteq V\backslash\{i,j,k\}$, such that $k\not\in\text{\emph{an}}_{G}(i,j,C)$, we have $i\mbox{$\>\perp\perp$ }j\,|\,C\cup\{k\}\Rightarrow i\mbox{$\>\perp\perp$ }j\,|\,C$.
@@
Denote $\text{sk}(G)$ to be the skeleton of graph $G$, formed by removing all arrowheads from edges in $G$. A v-configuration is a set of nodes $i,k,j$ such that $i$ and $j$ are connected to $k$, but $i$ and $j$ are not connected, and will be represented as $i\sim k\sim j$. A v-configuration oriented as $i\xrightarrow{}k\xleftarrow{}j$ is a collider, otherwise the v-configuration is a non-collider.
@@
To relate $\text{sk}(G)$ with distribution $P$, we define $\text{sk}(P)$ as follows:
Definition 4 ($\text{sk}(P)$).
Given a distribution $P$, the skeleton $\text{sk}(P)$ is the undirected graph with node set $V$, such that for all $i,j\in V$, the node $i$ is adjacent to $j$ if and only if there does not exist any $C\subset V\backslash\{i,j\}$ such that $i\mbox{$\>\perp\perp$ }j\,|\,C$.
Note that $\text{sk}(P)$ is the output of the skeleton building step of SGS/PC algorithm under faithfulness.
@@
Note that if $P$ is adjacency faithful wrt $G$, then $\text{sk}(P)=\text{sk}(G)$.
@@ 2.2 Natural Structure Learning Algorithms
Let $P$ be Markovian to the true causal graph $G_{0}$, then a causal learning algorithm aims to recover the graph $G_{0}$, up to the MEC, in which case we say that the algorithm is consistent. To relax the faithfulness assumption, Sad introduced natural structure learning algorithms.
@@
$P$ is V-stable if for all v-configurations $i\sim k\sim j$ in $\text{\emph{sk}}(P)$, and $C\subseteq V\backslash\{i,j,k\}$, the independencies $i\mbox{$\>\perp\perp$ }j\,|\,C$ and $i\mbox{$\>\perp\perp$ }j\,|\,C\cup\{k\}$ cannot both hold.
@@
This is a definition on $P$ itself, and is implied by the well-known singleton transitive axiom, under adjacency faithfulness.
Proposition 1 (Theorems 14 and 25 of [Sad]).
The graphs $G(P)$ and $G_{0}$ are Markov equivalent if the following holds:
@@
In [Sad], Condition 1 above is given in terms of converse pairwise Markovian instead of adjacency faithfulness, this is due to attempts in characterising the consistency conditions in terms of structural equation models (SEM). However, only the weaker adjacency faithfulness is needed and here we are focused on relaxing conditions.
By Example 21 in Sad, it can be seen that combined, these conditions are strictly weaker than restricted faithfulness, which is the weakest known consistency condition for SGS/PC [UhlSP].
Under the faithfulness assumption, constraint-based structure learning algorithms are natural. However, it is unclear whether these algorithms are still natural structure learning algorithms once the faithfulness assumption is relax, and no concrete algorithm is provided in [Sad]. Thus, without assuming faithfulness, we aim to provide a general concrete natural structure learning algorithm that relaxes the consistency conditions in Proposition 1.
As usual in constraint-based causal learning, we assume the availability of a conditional independence oracle—given a probability distribution $P$, we can determine with certainty whether conditional independence statements are true. In practice, conditional independence statements need to be estimated from the data using methods such as HSIC testing [HSIC], and is shown to be in general, a hard problem [hard].
@@ 3.1 Theory
Definition 8 (V-OUS and collider-stability).
A distribution $P$ is V-OUS and collider-stable wrt DAG $G$ if for all v-configuration $i\sim k\sim j$ in $G$:
@@
    (V-Ordered upward stability (V-OUS)). If $i\sim k\sim j$ is a non-collider, then for all $C\subseteq V\backslash\{i,j,k\}$, we have $i\mbox{$\>\perp\perp$ }j\,|\,C\Rightarrow i\mbox{$\>\perp\perp$ }j\,|\,C\cup\{k\}$.
@@
    (Collider-stability). If $i\xrightarrow[]{}k\xleftarrow{}j$, then there exists $C^{\prime}\subseteq V\backslash\{i,j,k\}$ such that $i\mbox{$\>\perp\perp$ }j\,|\,C^{\prime}$.
Collider-stability is related to orientation faithfulness, which states that the graph is faithful up to v-configurations in the graph [rams]. However collider-stability is much weaker, even than the Markovian assumption.
@@
V-OUS and collider-stability can be seen as local versions of ordered stabilities for the purposes of learning DAGs. In the case of DAGs, V-OUS can be seen as a relaxation of OUS since the implication $i\mbox{$\>\perp\perp$ }j\,|\,C\Rightarrow i\mbox{$\>\perp\perp$ }j\,|\,C\cup\{k\}$ in Definition 8 only has to hold for $i,j,k$ that are non-colliders in $G$. Likewise, collider-stable is implied by ODS and can be seen as a relaxation.
@@
Note that the above is the same with natural structure learning algorithms, just that one of the requirements is relaxed, namely Condition 2 in Definition 9. Thus, just like natural structure learning algorithms, all constraint-based algorithms that work under faithfulness are localised natural.
@@
A V-OUS and collider-stable orientation rule wrt $P$ is defined as an assignment of v-configurations $i\sim k\sim j$ in $\text{\emph{sk}}(P)$ into colliders and non-colliders as follows:
@@
    If for all $C$ such that $i\mbox{$\>\perp\perp$ }j\,|\,C$, we have $k\in C$, then assign $i\sim k\sim j$ to be a non-collider.
@@ A DAG $G$ is said to satisfy the V-OUS and collider-stable orientation rule wrt $P$, if $G$ satisfies:
    For all v-configurations $i\sim k\sim j$ in $G$, via the orientation rule in Definition 10,
@@
We have the following characterisation:
@@
We can now apply the V-OUS and collider-stable orientation rule wrt $P$ to assign the v-configurations in $\text{sk}(P)$. Note that the assignment may be incomplete, in the sense that some v-configurations may not satisfy either of the conditions in Definition 10 and are therefore unassigned. Modified V-stability is then defined as when this ambiguity can be resolved using the constraint that the graph is a DAG, as illustrated in Figure 1.
@@
A distribution $P$ is modified V-stable, if the v-configurations of DAGs that satisfy the V-OUS and collider-stable orientation rule wrt $P$ is unique.
Figure 1: Different assigned $\text{sk}(P)$. Label $\mathcal{N,C,U}$ denotes the v-configuration that is assigned to be a non-collider, collider and unassigned, respectively. In a), the unassigned v-configuration $4\sim 3\sim 2$ is constrained to be a collider due to acylicity of a DAG. In b), from lack of bidirectedness in a DAG, unassigned v-configuration $1\sim 2\sim 3$ is constrained to be a non-collider. In c), since there exists DAGs such that v-configuration $2\sim 3\sim 4$ can be either a collider or a non-collider, this ambiguity cannot be resolved.
$P$ is modified V-stable if, after orienting $\mathrm{sk}(P)$ via the V-OUS and collider-stable orientation rule wrt $P$, the unassigned v-configurations in $\mathrm{sk}(P)$ can be resolved using DAG constraints, as in a) and b).
Under V-stability, all v-configurations in $\text{sk}(P)$ must satisfy either of the conditions in Definition 10, leaving no v-configurations in $\text{sk}(P)$ unassigned. Thus, V-stability implies modified V-stability.
@@
We have the following, for the true causal graph $G_{0}$: